OpenStudy (anonymous):

What does "e" stand for in math? What's it used for? My answer: exponential-function base is the transcendental number e , which is equal to approximately 2.71828

OpenStudy (anonymous):

but what is it used for?? and what does "e" stand for?

OpenStudy (phi):

it is used to compute continuously compounded interest it is used for imaginary numbers it is used in representing signals (e.g. radio waves or sound waves)

OpenStudy (anonymous):

what does it stand for?

OpenStudy (anonymous):

There are many mathematical definitions of e \[\lim_{n \rightarrow \infty} (1+\frac{1}{n})^n\] \[e^{i\theta}=\cos(x)+i\sin(x)\] \[e^{i\theta}+1=0\] \[\frac{d}{dx}e^x=e^x\] \[\int e^xdx=x^x +c\] But one relation to rule them, one relation to find them all... \[e^x=1+x+\frac{1}{2!}x^2+\frac{1}{3!}x^3+\frac{1}{4!}x^4+...\] OR \[e^x=\sum_{k=0}^{\infty}\frac{x^k}{k!}\] It is possible to prove the all previous statements when we define e^x as above.

OpenStudy (anonymous):

huh?!!!!!!!! thats so confusing

OpenStudy (anonymous):

Which part is confusing?

OpenStudy (anonymous):

And why?

OpenStudy (anonymous):

Take a look here for more explanation: http://betterexplained.com/articles/an-intuitive-guide-to-exponential-functions-e/ betterexplained.com/articles/intuitive-understanding-of-eulers-formula/ http://en.wikipedia.org/wiki/E_%28mathematical_constant%29 And knowing the Taylor series is a good idea too.